11. Suppose that a perfectly competitive market is described by the following demand and supply equations QD 10-P What is the equilibrium price in this market? A. $2 $3 C.S4 D.S5 E. S6 12. Suppose that a perfectly competitive market is described by the following demand and supply equations Q-10-P What is the value of total social surplus (consumer and producer surplus) in this market in equilibrium? A. 9 B. $18 C. S27 D. S36 E. S54I found the answer to number 11, which is C $4, however I am not sure how to find the answer to number 12.
Answer #1

Answer 12

The correct answer is (C) $27

Producer Surplus is the benefit Producer receives by selling at a price more than what they are willing to sell. Mathematically Producer Surplus is the area above Supply curve and below equilibrium Price.

Consumer Surplus is the benefit Consumer receives by paying a price Less than what they are willing to pay. Mathematically Consumer Surplus is the area below demand curve and above equilibrium Price.

As You have Calculated above equilibrium Price = 4. Hence equilibrium quantity = Q = 10 - 4 = 6

Dena I D

We can see from above figure and above definitions that, Consumer Surplus = A

We can also see from above figure and above definitions that, Producer Surplus = B

Hence Total Surplus = A + B

Area of a triangle = 0.5*base*height

=> Total Surplus = (1/2)*(10 - 1)*6 = 27

Hence Total Surplus = $27

Hence the correct answer is (C) $27